Mr. Hernandez had to interview a total of 63
households in his assignment. He has already
finished 28. What percentage of the households
in his assignments has he finished? (Round
your answer to the nearest tenth of a percent.)

Answer by richwmiller(17219)

28/63=.444=44.4% which makes sense since 28/63 is almost 30/60=1/2